Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured iComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Req 3 Req 4 Req 5 Req 6A Req 6B Compute (a) laDue to an increase in labor rates, the company estimates that next years variable expenses will increase by $3.00 per ball.Refer to the data in Required (2). If the expected change in variable expenses takes place, how many balls will have to be soRefer again to the data in Required (2). The president feels that the company must raise the selling price of its basketballsRefer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plantIf the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $128,000, asAssume the new plant is built and that next year the company manufactures and sells 34,000 balls (the same number as sold las
Answer #1
Req 1
CM Ratio 40.00%
Unit Sales to break even 21200
Degree of operating leverate 1.60
CM ratio = contribution /sales
340,000/850,000
40.00%
BEP(units) = fixed cost/contribution margin per unit
212000/10
21200
Degree of operating leverage = contribution/net income
340000/212000
1.60
Req 2 CM Ratio 28.00%
Unit Sales to break even 30286
CM ratio = contribution /sales
7./25
28.00%
BEP(units) = fixed cost/contribution margin per unit
212000/7
30286
Req 3
Number of balls 48571
BEP(units) =( fixed cost+ target income)/contribution margin per unit
(212000+128000)/7
48571
Req 4 selling price 30.00
CM ratio = 40%
selling price per unit be x
variable cost per unit is 18
so selling price should be =
40%              = (x-18)/x
.40x                  = (x -18)
x                         =18/.6
x                         = 30.00
Req 5 Selling price per unit 25
New variable cost (15*60%) 9
Contribution per unit 16
fixed expense   = 212000*200%= 424000
contribution margin ratio 64.00%
unit sales to break-even 26500 balls
Req 6A number of balls 34500 balls
(424000+128000)/16
Req6B Contribution income statement
Sales (34000*25) 850000
Variable expenses (34000*9) 306000
Contribution margin 544000
Fixed expenses 424,000
Net operating income 120,000
Degree of operating leverage 4.53
(contribution margin/net income)